OpenStudy (anonymous):

how to integrate 2x+7/(2x+1)(x+2) with limits 7,0 into ln50

OpenStudy (anonymous):

multiply and divide the second term by ((2x+1)-2(x+2))

ganeshie8 (ganeshie8):

\(\large \int \limits_0^7 \frac{2x+7}{(2x+1)(x+2)} dx\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Yup!

ganeshie8 (ganeshie8):

familiar with partial fractions ?

OpenStudy (anonymous):

You'll need to use partial fractions (C4 Chapter 1) to simplify, then integrate.

OpenStudy (anonymous):

Kind of. I got 4/2x+1 -1/x+2

ganeshie8 (ganeshie8):

\(\large \frac{2x+7}{(2x+1)(x+2)} = \frac{A}{2x+1} + \frac{B}{x+2}\)

OpenStudy (anonymous):

Yup, I got that.

ganeshie8 (ganeshie8):

\(\large \frac{2x+7}{(2x+1)(x+2)} = \frac{4}{2x+1} + \frac{-1}{x+2}\)

ganeshie8 (ganeshie8):

looks good :)

ganeshie8 (ganeshie8):

\(\large \int \limits_0^7 \frac{2x+7}{(2x+1)(x+2)}~dx = \int \limits_0^7 \frac{4}{2x+1} + \frac{-1}{x+2}~dx\)

ganeshie8 (ganeshie8):

and why is that easy to integrate now ?

ganeshie8 (ganeshie8):

it is easy cuz we knw : \(\int \frac{1}{x}~ dx = \ln x\)

ganeshie8 (ganeshie8):

you just need to fix the constants okay

ganeshie8 (ganeshie8):

\(\large \int \limits_0^7 \frac{2x+7}{(2x+1)(x+2)}~dx = \int \limits_0^7 \frac{4}{2x+1} + \frac{-1}{x+2}~dx\) \(\large ~~~~~~~~~~~~~~~~~~= \int \limits_0^7 \frac{4}{2x+1}~dx + \int \limits_0^7 \frac{-1}{x+2}~dx\) \(\large ~~~~~~~~~~~~~~~~~~= 4\int \limits_0^7 \frac{1}{2x+1}~dx - \int \limits_0^7 \frac{1}{x+2}~dx\) \(\large ~~~~~~~~~~~~~~~~~~= 4 \frac{\ln (2x+1)}{2}\Big|_0^7 - \frac{1}{x+2}\Big|_0^7\)

ganeshie8 (ganeshie8):

see if that makes sense so far, next, u just need to take the limits and simplify

OpenStudy (anonymous):

Thanks hehe!

ganeshie8 (ganeshie8):

u wlc :) when u take the limits it does simplify to ln50

ganeshie8 (ganeshie8):

good luck !

OpenStudy (anonymous):

I got ln25 ><

ganeshie8 (ganeshie8):

check ur calculation again

OpenStudy (anonymous):

I took (15^2)/2 then /9

ganeshie8 (ganeshie8):

\(~~~~~~~~~~~~~~~~~~= 4 \frac{\ln (2x+1)}{2}\Big|_0^7 - \ln(x+2)\Big|_0^7\) \(~~~~~~~~~~~~~~~~~~= 2 \ln (2x+1)\Big|_0^7 - \ln(x+2)\Big|_0^7\) \(~~~~~~~~~~~~~~~~~~= 2 \left[\ln (2*7 + 1) - \ln(2*0+1)\right] - \left[\ln(7+2) - \ln(0+2)\right]\) \(~~~~~~~~~~~~~~~~~~= 2 \left[\ln (15) - \ln(1)\right] - \left[\ln(9) - \ln(2)\right]\) \(~~~~~~~~~~~~~~~~~~= 2 \left[\ln (15) \right] - \left[\ln(9) - \ln(2)\right]\) \(~~~~~~~~~~~~~~~~~~= \ln (15^2) - \ln(9) + \ln(2)\) \(~~~~~~~~~~~~~~~~~~= \ln (\frac{15^2\times 2}{9}) \) \(~~~~~~~~~~~~~~~~~~= \ln (50) \)

ganeshie8 (ganeshie8):

see if that looks okay...

OpenStudy (anonymous):

Omg, thanks!!!! I forgot to separate the two integrals and forgot to sub in 0 ><

ganeshie8 (ganeshie8):

np :)